OpenStudy (erikaxx):

QUESTION: Find the area of the first arch of the curve y = e^(-x) sin x TOPIC: PLANE AREAS IN RECTANGULAR COORDINATES IN CALCULUS

OpenStudy (erikaxx):

\[y = e^{-x} sinx\]

OpenStudy (erikaxx):

@kittiwitti1

OpenStudy (erikaxx):

@NoelGreco

OpenStudy (kittiwitti1):

I can't help, sorry. I'm also working on a problem of my own at the minute -- @ganeshie8 ?

OpenStudy (noelgreco):

1. Find the zeros. Remember that the "e" factor is always ppositive 2. Integrate by parts twice 3. Evaluate

OpenStudy (erikaxx):

are the limits is from 0 to 0?

OpenStudy (erikaxx):

can you explain it one by one so i can understand please

OpenStudy (irishboy123):

you want \(\int\limits_0^{\pi / 2} \sin x \; e^{-x} \; dx\), right?! so in an IBP the \(\sin\) will flip back and forward from \(\sin \to \cos \to \sin \to \cos \dots\), and the exponential just changes from \(+ \to - \to + \to - \dots\) so you can pretty much do what you fancy, quite a luxury really :p

OpenStudy (erikaxx):

so by ibp i will get

OpenStudy (erikaxx):

so by ibp i will get \[\int\limits e ^{-x} sinx dx \]

OpenStudy (erikaxx):

\[u = sinx, du = \cos x dx ,dv = \int\limits e ^{-x}dx , v =e ^{x}\]

OpenStudy (erikaxx):

\[=e ^{-x}sinx dx - \int\limits e ^{-x}\cos xdx\]

OpenStudy (erikaxx):

\[u^`= cosx, du^` = `sinx dx , dv = e ^{-x} \cos xdx , v = e^^{-x}\]

OpenStudy (erikaxx):

then i will get a equation of \[1/2 e ^{-x} (sinx - \cos x)+ c \]

OpenStudy (erikaxx):

@SolomonZelman

OpenStudy (solomonzelman):

As you were directed, an integration by parts twice will allow you to solve the integral. You will integrate it without completely integrating it. Want an example of such?

OpenStudy (erikaxx):

yes

OpenStudy (erikaxx):

thats my answer

OpenStudy (erikaxx):

is it right?

OpenStudy (solomonzelman):

Suppose you have the following integral: \(\color{#000000}{ \displaystyle \int e^x\cos x~dx }\) Integrate by parts. Set, \(\color{#000000}{ \displaystyle u=\cos x \quad\quad\quad dv=e^x~dx }\) And from that setup you have, \(\color{#000000}{ \displaystyle du=-\sin x~dx \quad\quad\quad v=e^x }\) then, you get: \(\color{#000000}{ \displaystyle \int e^x\cos x~dx=e^x\cos x-\left(\int(e^x)(-\sin x~dx)\right) }\) \(\color{#000000}{ \displaystyle \int e^x\cos x~dx=e^x\cos x+\int e^x\sin x~dx }\) Then, you will apply integration by parts to the second integral with e^x•sin(x): \(\color{#000000}{ \displaystyle \int e^x\sin x~dx=? }\) Set, \(\color{#000000}{ \displaystyle u=\sin x \quad\quad\quad dv=e^x~dx }\) from there, \(\color{#000000}{ \displaystyle du=\cos x~dx \quad\quad\quad v=e^x }\) \(\color{#000000}{ \displaystyle \int e^x\sin x~dx=e^x\sin x-\int e^x\cos x~dx }\) Going back to the previous integral you have: \(\color{#000000}{ \displaystyle \int e^x\cos x~dx=e^x\cos x+\int e^x\sin x~dx }\) \(\color{#000000}{ \displaystyle \int e^x\cos x~dx=e^x\cos x+\left(e^x\sin x-\int e^x\cos x~dx\right) }\) Then, I will rely on your algebra: \(\color{#000000}{ \displaystyle \int e^x\cos x~dx+\int e^x\cos x~dx=e^x\cos x+e^x\sin x }\) \(\color{#000000}{ \displaystyle 2\int e^x\cos x~dx=e^x\cos x+e^x\sin x }\) \(\color{#000000}{ \displaystyle \int e^x\cos x~dx=\frac{e^x\cos x+e^x\sin x}{2} }\)

OpenStudy (solomonzelman):

I think no right. v = \(\color{red}{-}\)e^(-x).

OpenStudy (solomonzelman):

since the integral of e^(-x) is -e^(-x)

OpenStudy (erikaxx):

yes, same thing that i integrate, did u see the jpeg that i attached?

OpenStudy (solomonzelman):

Yes, I did, and that is exactly I am saying wha I said just now.

OpenStudy (solomonzelman):

exactly **why** I am saying ...

OpenStudy (solomonzelman):

so in your first integration by parts, you would be actually adding the integral, not subtracting (because there are two negatives, (1) you subtract the integral when integrating by parts, (2) the -e^(-x), which is the integral of dv. So --=+)

OpenStudy (erikaxx):

why? the formula of ibp is uv - integral of vdu

OpenStudy (solomonzelman):

yes, and asides from the negative in the formula, you have another negative coming from -e^(-x).

OpenStudy (solomonzelman):

So you have two negatives, and that is plus.

OpenStudy (solomonzelman):

You error is that the integral of e^(-x) is -e^(-x), and you missed that negative.

OpenStudy (solomonzelman):

you missed the minus in front of the e^(-x).

OpenStudy (erikaxx):

but the negative is in exponent so nothing will do with that? @SolomonZelman

OpenStudy (solomonzelman):

Ok, what is the integral of e^(-x)?

OpenStudy (erikaxx):

1/ dx

OpenStudy (solomonzelman):

That is the answer to my question?

OpenStudy (erikaxx):

no, -e^(-x)

OpenStudy (solomonzelman):

Yes, exactly.

OpenStudy (erikaxx):

can you help me understand it by steps,

OpenStudy (solomonzelman):

So, you have: \(\color{#000000}{ \displaystyle \int e^{-x}\sin x dx }\) Then you set, \(\color{#000000}{ \displaystyle u=\sin x }\), and so \(\color{#000000}{ \displaystyle du=\cos x~dx }\) \(\color{#000000}{ \displaystyle dv=e^{-x}~dx }\), and so \(\color{#000000}{ \displaystyle v=-e^{-x} }\) The formula is: \(\color{#000000}{ \displaystyle \int u~dv=uv-\int v~du }\) And here you get: \(\color{#000000}{ \displaystyle \int(\sin x)( e^{-x} ~dx)=\sin x(-e^{-x})-\int (-e^{-x})~(\cos x~dx) }\)

OpenStudy (solomonzelman):

So, you should have had the following integral: \(\color{#000000}{ \displaystyle \int(\sin x)( e^{-x} ~dx)=-e^{-x}\sin x+\int e^{-x}\cos x~dx }\)

OpenStudy (erikaxx):

ah yes, so because we need 2 integration i will next integrate the \[\int\limits e ^{-x} \cos x\]

OpenStudy (solomonzelman):

Yes, indeed!

OpenStudy (solomonzelman):

So, now, YOU need to integrate this integral....

OpenStudy (solomonzelman):

well, you should be consistent. So, you will need to be integrating e^(-x) again. (So, you can't choose cosine as dv, because then you are just going to return to your initial integral.)

OpenStudy (solomonzelman):

So you must set dv=e^(-x). and thus u=cos(x).

OpenStudy (solomonzelman):

dv=e^(-x) dx I meant.

OpenStudy (erikaxx):

here

OpenStudy (erikaxx):

is that right

OpenStudy (erikaxx):

@SolomonZelman

OpenStudy (solomonzelman):

yes, I am there... Please check your sines.

OpenStudy (solomonzelman):

I mean signs

OpenStudy (erikaxx):

oops its should b positive at \[2 \int\limits e ^{-x} \sin x dx = -e ^{-x} sinx + e ^{-x} \cos x\]

OpenStudy (erikaxx):

so it would five a answer of \[\int\limits e ^{-x} \sin x dx = -e ^{-x} (sinx + cosx)/ 2 + c\]

OpenStudy (solomonzelman):

Yes, very good! \(\color{#000000}{ \displaystyle \int(\sin x)( e^{-x} ~dx)=\frac{-e^{-x}}{2}(\sin x+ \cos x)+C }\)

OpenStudy (erikaxx):

yes, it is will be divide by 2?

OpenStudy (solomonzelman):

?

OpenStudy (solomonzelman):

elaborate

OpenStudy (solomonzelman):

I just didn't quite understand what you asked.

OpenStudy (solomonzelman):

In any case, the first arc, from 0 to π/2. That is just plugging in the limits of integration and preforming the algebraic task. I think you can do this without me. Good luck with your maths:)

OpenStudy (erikaxx):

i mean the whole equation of sinx + cos x will be divided by 2

OpenStudy (erikaxx):

no i can`t still @SolomonZelman

OpenStudy (solomonzelman):

Yes, it will be. (You can use my answer, where I essentially factored the exponential with the 1/2)

OpenStudy (solomonzelman):

We had twice the integral on the left, so to solve for the integral we needed to divide by 2, correct ?

OpenStudy (erikaxx):

so what about the first arc thingy in the question?

OpenStudy (erikaxx):

yep

OpenStudy (solomonzelman):

Well, that is this integral from 0 to \(\pi/2\).

OpenStudy (solomonzelman):

So you just plug in the limits of integration.... that on its own is just algebra.

OpenStudy (erikaxx):

i dont know how to compute e

OpenStudy (solomonzelman):

Well, show me your work to the point where you are stock, please.

OpenStudy (erikaxx):

here

OpenStudy (erikaxx):

i dont know how to compute in calcu it gives me a decimal answer

OpenStudy (solomonzelman):

You can leave the answer in exact form. You can be very fancy and say: \(e^{-\pi/2}=i^i\) You can approximate it as a decimal. it all depends on what your teacher would want.

OpenStudy (solomonzelman):

I assume that you know cosine and sine of π and of π/2, tho. Can I assume so?

OpenStudy (solomonzelman):

I would say that just getting rid of negative exponents, but to leave the answer in exact-simplified form is the best thing to do.

OpenStudy (erikaxx):

but my teacher want a not decimal answer

OpenStudy (solomonzelman):

So, then let's get the exact answer.

OpenStudy (solomonzelman):

I am lagging this whole time. Sorry

OpenStudy (erikaxx):

okay so the answer is 0?

OpenStudy (solomonzelman):

It is not

OpenStudy (erikaxx):

i got 0 because pi/2 is equivalent to 180 degree, so by replacing it 0-1+0+1 = 0

OpenStudy (solomonzelman):

You are missing something.

OpenStudy (solomonzelman):

the first (0-1) would be multiplied by e^(-π/2), and the second (0+1) would be multiplied by 1/2, since e^(0)=1.

OpenStudy (solomonzelman):

At first, lets agree that we are evaluating the following: \(\color{#000000}{ \displaystyle -\frac{e^{-\pi/2}}{2} (\sin{\tiny~} \pi/2+\cos{\tiny~}\pi/2) -\left\{-\frac{e^{-0}}{2} (\sin{\tiny~} 0+\cos{\tiny~}0) \right\} }\) (result of plugging the limits of integration.

OpenStudy (erikaxx):

i cant compute the e^ (-pi/2) caus ethe calcu says 0.20

OpenStudy (solomonzelman):

Your teacher wants a non-decimal answer, so why not give him/her the exact answer.

OpenStudy (solomonzelman):

We will enchance the answer by applying some algebra, but please try to follow with me as far as you can.

OpenStudy (solomonzelman):

We are really evaluating the following: \(\color{#000000}{ \displaystyle -\frac{e^{-\pi/2}}{2} (\sin{\tiny~} \pi/2+\cos{\tiny~}\pi/2) +\frac{e^{-0}}{2} (\sin{\tiny~} 0+\cos{\tiny~}0) }\) do you understand why?

OpenStudy (erikaxx):

okay i will follow you

OpenStudy (erikaxx):

so 2 will be cancel out and e^pi will remain

OpenStudy (solomonzelman):

Can you tell me: \(e^{-0}=e^0=?\)

OpenStudy (erikaxx):

is equals to 1

OpenStudy (solomonzelman):

So, let's re-write the expression: \(\color{#000000}{ \displaystyle -\frac{e^{-\pi/2}}{2} (\sin{\tiny~} \pi/2+\cos{\tiny~}\pi/2) +\frac{1}{2} (\sin{\tiny~} 0+\cos{\tiny~}0) }\) YOu can evaluate step by step, if quicker wrking results in small errors.

OpenStudy (erikaxx):

okay then what to do with the first e how will going to eliminate it

OpenStudy (solomonzelman):

And as you said: \(\sin 0=\cos \pi/2 =0\) \(\sin \pi/2=\cos 0 =1\) So you would then end up with: \(\color{#000000}{ \displaystyle -\frac{e^{-\pi/2}}{2} (1+0) +\frac{1}{2} (0+1) }\) \(\color{#000000}{ \displaystyle -\frac{e^{-\pi/2}}{2} +\frac{1}{2} }\) \(\color{#000000}{ \displaystyle \frac{1}{2}-\frac{e^{-\pi/2}}{2} }\)

OpenStudy (solomonzelman):

And then what you so much desired. But, the bad news is that you can't eliminate it, really. THo, you could re-write it without negative exponent.